OpenStudy (anonymous):

{3,10,29,66,127,218} write in set builder form.

mathslover (mathslover):

@Deepakmalhotra , first of all welcome to openstudy . Deepak, can you identify any pattern in them?

mathslover (mathslover):

Hint : \(1^3 + 2 = 3\) and \(2^3 + 2 = 8 \) \(...\) \(6^3 + 2 = 216 + 2\) So it is the form : \(n^3 + 2\) (talking about the series : \(3,10,29,66,127,218\) ) Can you write the given data : \(\{3,10,29,66,127,218\}\) in set - builder form now?

mathslover (mathslover):

Note some conditions also : \(x\in \mathbb{N}\) \(x^3 + 2 \space \textbf{is the pattern followed}\)

mathslover (mathslover):

It seems you are having problem in replying (we talked in message box) check the attachment.

OpenStudy (anonymous):

yaa yaa

mathslover (mathslover):

good! So can you tell me what will be the set builder form of the given series?

OpenStudy (anonymous):

i also dont...............know....

mathslover (mathslover):

Can you write set builder form for : \(\{ 2,4\} \) ?

OpenStudy (anonymous):

ok ...

OpenStudy (anonymous):

A={x:x is even natural no. and 1<x<5}

mathslover (mathslover):

Now can u write for : \(\{2,5,10,17\}\) ?

OpenStudy (anonymous):

sorry i will write as soon as light will come ...eletricity has gone....

mathslover (mathslover):

No problem, I will not be present here : \(\{x:x^3+2 , x\in \mathbb{N}\}\) [this is the answer as , the numbers are following the pattern of x^3 + 2] and the limitation is : [x belongs to N(Natural Numbers)]

Parth (parthkohli):

Erm no, it's actually...\[\{x^3 +2 : x \in \mathbb{N}, x \le 6\}\]@mathslover It's not the set of \(x\)'s and it's not an infinite set. It's the set of \(x^3 + 2\) such that \(x\) belongs to natural numbers and is less than or equal to 6. :-|

Parth (parthkohli):

@mathslover

mathslover (mathslover):

\(\large{\{y:x^3 + 2 = y, x \in \mathbb{N} , x\le 6\}}\) This is better, and Parth it can also be used in finite set. why did you include " its not an infinite set' ?

Parth (parthkohli):

@mathslover because the question doesn't say that it's an infinite set.